OpenStudy (anonymous):

a horizontal force of 50N is needed to keep a 25kg block moving with a constant velocity of magnitude 5m/s across a horizontal table. what is the coefficient of kinetic friction?

OpenStudy (anonymous):

do you knw the friction?

OpenStudy (anonymous):

nvm the friction.... so the thing is... Fapplied-Ffriction = Fnet and Fnet = 0 because speed is constant... so, Fapplied = 50 therefore, u = coefficient of friction 50-uFnormal = Fnet Fnormal is mg = 245 N so, Fnet = Fapplied - uFnormal 0 = 50 -u245 u245 = 50 u = 50/245 u = 0.20 so, my dear frnd, coefficient of friction is 0.20

OpenStudy (anonymous):

thanks alot man

OpenStudy (anonymous):

no problem :D